# AttriBoT: A Bag of Tricks for Efficiently Approximating Leave-One-Out Context Attribution

### v2

@@ Abstract
The influence of contextual input on the behavior of large language models (LLMs) has prompted the development of context attribution methods that aim to quantify each context span’s effect on an LLM’s generations. The leave-one-out (LOO) error, which measures the change in the likelihood of the LLM’s response when a given span of the context is removed, provides a principled way to perform context attribution, but can be prohibitively expensive to compute for large models. In this work, we introduce AttriBoT, a series of novel techniques for efficiently computing an approximation of the LOO error for context attribution. Specifically, AttriBoT uses cached activations to avoid redundant operations, performs hierarchical attribution to reduce computation, and emulates the behavior of large target models with smaller proxy models. Taken together, AttriBoT can provide a >300$\times$ speedup while remaining more faithful to a target model’s LOO error than prior context attribution methods. This stark increase in performance makes computing context attributions for a given response 30×30$\times$30 × faster than generating the response itself, empowering real-world applications that require computing attributions at scale. We release a user-friendly and efficient implementation of AttriBoT to enable efficient LLM interpretability as well as encourage future development of efficient context attribution methods 222 https://github.com/r-three/AttriBoT.
@@ 1 Introduction
A natural approach for context attribution is to remove a span of text from the context and measure the ensuing change in the likelihood of the model’s original response. This notion of importance, known as the Leave-One-Out (LOO) error, is a common idea used in training data attribution (Koh & Liang, 2020), data valuation (Choe et al., 2024), feature attribution (Li et al., 2017), and recently has been explored for context attribution (Cohen-Wang et al., 2024). While the LOO error produces meaningful and interpretable context attribution scores, it is often viewed as impractical due to the need to perform an independent forward pass to score each text span in the context. This is particularly problematic for modern LLMs whose forward passes are computationally expensive and realistic use cases often involve long contexts with many text spans (e.g., sentences or paragraphs) to score.
@@
These findings naturally lend themselves to efficient key-value caching schemes, novel hierarchical approaches that prune low-information context chunks before performing attribution, and methods that leverage small proxy LMs to approximate the LOO attributions of a larger target model. We develop and explore the practical application of these schemes to develop an aggregate system called AttriBoT (a Bag of Tricks for efficient context attribution). When evaluating AttriBoT in the open-book question answering setting, where a model is presented a question with one or more related documents provided in its context, we find that our methods significantly reduce the cost of computing LOO attributions – at times by >300×300$\times$300 × – while remaining more faithful to the original model’s LOO error than a wide range of baselines. In addition, the cascade of approaches underlying AttriBoT can be naturally composed to attain a Pareto-optimal trade-off between efficiency and accuracy over efficiencies ranging multiple orders of magnitude.
@@ 6 Conclusion
In this paper, we introduced AttriBoT, a Bag of Tricks for efficiently attributing an LLM’s predictions to spans of text in its input context. AttiBoT leverages a series of novel techniques for approximating the attributions of a large target model, including reusing computations through key-value caching, performing hierarchical attributions to reduce the number of forward passes, and leveraging a smaller proxy model whose attributions reliably approximate those of the target model. Taken together, AttriBoT can provide a >300$\times$ speedup, making attributing a response 30×30$\times$30 × more efficient than even generating a response in realistic settings, while remaining more faithful to the target model’s attributions than prior methods for efficient context attribution. In addition, the components of AttriBoT can be included, excluded, and tuned to produce a Pareto-optimal trade-off between faithfulness and speed. While our experimental results primarily focus on the general setting of open-book question answering, we anticipate AttriBoT will also be useful for detecting malicious prompts and model hallucinations. We hope that our efficient and easy-to-use implementation of AttriBoT2 ensures that it has real-world impact and also enables future work on efficient context attribution.

### v3

@@ 1 Introduction
A natural approach for context attribution is to remove a span of text from the context and measure the ensuing change in the likelihood of the model’s original response. This notion of importance, known as the Leave-One-Out (LOO) error, is a common idea used in training data attribution (Koh & Liang, 2020), data valuation (Choe et al., 2024), feature attribution (Li et al., 2017), and recently for context attribution (Cohen-Wang et al., 2024). While the LOO error produces meaningful and interpretable context attribution scores, it is often viewed as impractical due to the need to perform an independent forward pass to score each text span in the context. This is particularly problematic for modern LLMs whose forward passes are computationally expensive and realistic use cases often involve long contexts with many text spans (e.g., sentences or paragraphs) to score.
@@
[…]
@@ 6 Conclusion
[…]